I need help solving the following problems using the square root property:
3(x -5)^2 - 7 = - 4
(4x -2)^2 = -16

Answer by checkley75(3666)

3(X-5)^2-7=-4
3(X-5)^2=7-4
3(X-5)^2=3
(X-5)^2=3/3
(X-5)^2=1 NOW TAKING THE SQUARE ROOT OF BOTH SIDES WE GET:
X-5=1
X=1+5
X=6 ANSWER.
-----------------------------------------------
(4X-2)^2=-16 AGAIN SQUARING BOTH SIDES WE GET:
4X-2=256
4X=256+2
4X=258
X=258/4
X=64.5 ANSWER.